I need to identify the center, foci, and graph the following
(x+1)^2/36+y^2/49=1

Answer by lwsshak3(11628)

I need to identify the center, foci, and graph the following
(x+1)^2/36+y^2/49=1
This is an equation of an ellipse with vertical major axis of the standard form:
(x-h)^2/b^2+(y-k)^2/a^2=1
For given equation:
center: (-1,0)
a^2=49
a=√49=7
b^2=36
b=√36=6
c^2=a^2-b^2=49-36=15
c=√15≈3.87
Foci:(-1,0±c)=(-1,0±√15)=(-1,0±3.87)=(-1,-3.87) and (-1,3.87)
see graph below:
y=±(49-49(x+1)^2/36)^.5